To properly apply some fertilizer to your home’s lawn you must spread 0.51 pounds of fertilizer for every square foot of your home’s lawn. Your lawn is 39.8 meters long and 20.2 meters wide. How many pounds of fertilizer do you need to buy? Note: 1 meter = 3.28 feet
1
Expert's answer
2016-01-26T09:50:20-0500
First of all the dimensions of lawn should be expressed as feet:
L = 39.8 m = 39.8×3.28 ft = 130.544 ft
W = 20.2 m = 20.2×3.28 ft = 66.256 ft
Then the square of the lawn is:
S = L×W = 130.544 ft × 66.256 ft = 8649.32 ft2
If 1 ft2 corresponds 0.51 pounds, the amount of fertilizer which we need to buy is defined:
M = 0.51 × S = 0.51 × 8649.32 lb = 4411.16 lb